OpenStudy (anonymous):

5x + 6y = 10 solve for y

OpenStudy (haichi):

Are youmissing any information

OpenStudy (anonymous):

no

OpenStudy (haichi):

you * ?*

OpenStudy (haichi):

hrmm.....

OpenStudy (anonymous):

Subtract 5 to both sides. Start from there.

OpenStudy (anonymous):

its a literal equation guys

OpenStudy (anonymous):

5x + 6y = 10 6y = 10-5x y = (10-5x)/6 y = 10/6-(5/6)*x y = 5/3 - (5/6)*x = 5/3- 5x/6

OpenStudy (anonymous):

\[5x+6y=10\] \[6y=5\] \[y=\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

(10-5x)/6 was CORRECT THANK U ALL